if two boys pull a rope "eastward" with forces of 40n and 30n, what are the magnitude and direction of the resultant? of the equilibrant?
1
Expert's answer
2018-10-05T14:39:09-0400
The magnitude:
F_net=F_1+F_2=40+30=70 N.
The direction is eastward.